# Quasi-randomization tests for network interference

### v2

@@ 4.2.2 Test for exponential random graph model
Many sufficient statistics used in real-life social network modeling follow this property. As discussed, some graph statistics include the number of triangles, degree sequence, and number of stars of size two, and they are permutation invariant. We model the data-generating process as an exponential random graph with permutation invariance property and present a permutation testing procedure.
@@
    $[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\;|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
@@
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a cluster randomized experiment. We assume that $G_{obs}\sim P_{G}$ where $G$ is a permutation-invariant exponential random graph model. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 3 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
@@ 4.3 Test statistic
It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty. For example, if no control unit exists with all neighbors in control or the number of such units is few, we deal with either a definition or power issue. Thus, we propose a generalization of the above-stated test statistic for robustness. We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units. Let $T_{quant_{c}}(G|Z_{obs})=$
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}$ (6)
@@
We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes. This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)). One can also build test statistics based on models for interference structures. For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993). We also consider more test statistics, given that the choice is not unique and different test statistics would differ in their power. We take the test statistic proposed in Bond et al. (2012), which takes the difference between the average over all the edges where the neighbor is a treated or control unit.
@@ 5.1 Data generating mechanism
We now present the data-generating process considered in the simulations. Our potential outcome function is linearly separable in direct treatment effect and spillover effect. The spillover mechanism is a function of the unit’s neighbor. We propose the mechanism to be contingent upon the proportion of treated neighbors.
@@
We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated. Without loss of generality, we take the mean of the normal distribution to be zero, but it can be arbitrarily chosen. We first present the setup for a completely randomized experiment. The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and assigning them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). We state that $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have $599$ total units and $300$ treated units in the model in one case. We study the setup for two generating processes: small-world networks and the stochastic block model. Both models exhibit a high degree of clustering, characterized by the clustering coefficient. The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible. Consider a regular graph where all the nodes have the same cardinality of connections. Let us denote the number of connections a unit has in a regular graph by $K$. Such graphs have a high clustering coefficient, but the distance between two nodes in such graphs is also high.
Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment.
$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$ $T_{elc}$
Small world network 0 0 0.045 0.049 0.049 0.047 0.050
4 0 0.054 0.046 0.056 0.045 0.046
0 0.4 0.187 0.148 0.191 0.162 0.112111We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
4 0.4 0.074 0.151 0.064 0.147 0.060222We obtain the same rejection rate in our replication as reported in Athey et al. (2015).
Stochastic block model 0 0 0.045 0.062 0.050 0.050 0.054
4 0 0.050 0.055 0.047 0.052 0.052
0 0.4 0.191 0.217 0.189 0.216 0.065
4 0.4 0.064 0.220 0.063 0.217 0.048
The average distance between two nodes in real-life networks is low (Amaral et al. (2000)). To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$. This parametrization characterizes a small world network. The higher the rewiring probability, the lower the average path length. Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model. Consider a graph of size $N$ with a partition of nodes. The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$. These parts represent clusters in the graph, and we can model a random graph using it by defining the probability of an edge between nodes within a part and between two parts. Hence, we obtain a probability (symmetric) matrix of size $K\times K$. Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements. The probability matrix is also called the preference matrix in the literature. We now define the setup for the simulation study:
@@
    Network: We first present results taking network distribution as a small-world network with the following parameters: K (initial connection of all units) = 10 and $p_{rw}$ (probability of rewiring) = 0.1. We then consider a stochastic block model with block sizes $\{50,100,40,110,299\}$. The preference matrix which specifies the probability of an edge between two groups is given below:
@@
The null distribution was generated by sampling 1000 graphs with the same degree sequence as the observed graph and isomorphic to the observed graph in a second option. The p-value is calculated as the sample proportion of times the imputed test statistic for the sampled graph realizations is more than the observed test statistic. We reject the null hypothesis if the p-value obtained is less than the chosen significance level of 0.05. We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null. We present this result, namely rejection rate, in the Table 1. For robustness, we perform the simulations for two network-generating processes. We obtain the rejection rates for the same setup for the procedure proposed by Athey et al. (2018). We implement random focal vertex selection and show our results for the Edge-Level-Contrast test statistic (Athey et al. (2018)). We label this test as ‘Conditional Focal Test’.
We then perform the simulations for cluster randomized experimental setup. We present the results in Table 2. The total number of units in the simulation is the same as before, that is, 599. We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Appendix A). We take epsilon to be equal to 3. This implies all the clusters contain units less than three hops away from each other, whereas the between-cluster distance is at least three hops for any two units. We take the probability of treating a cluster to be 0.5. We implemented a Bernoulli trial at the cluster level. Under Procedure 3, the null distribution is approximated by taking 1000 samples of graphs isomorphic to the original graph, such that the corresponding subgraphs on treated units and controlled units remain isomorphic. This is obtained by considering permutations where treated units can be mapped only to treated units (and thus, control units can only be mapped to control units). We replicated this setup 4000 times and calculated the proportion of times the proposed test rejected the null. We label Procedure 3 as ‘Block Isomorphism Permutation Test’ in Table 2. This was performed for two network-generating processes as before.
@@
Small world network 0 0 0.052 0.053
4 0 0.056 0.051
0 0.4 0.429 0.352
4 0.4 0.429 0.348
Stochastic block model 0 0 0.051 0.057
4 0 0.048 0.046
0 0.4 0.225 0.237
4 0.4 0.227 0.242
@@ 5.2 Comparison with related work
We see that our proposed methods have type 1 error controlled appropriately. This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2. When $\tau_{spill}$ equals 0.4, our method performs with competitive power compared to existing methods in the literature (Athey et al. (2018)). We highlight the $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method significantly improves. The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al. (2015)). We see in Table 1 that our proposed method boosts power when the clustering of the network-generating process increases, as with the stochastic block model. This contrasts with the method proposed by Athey et al. (2018), where the method does not retain power. In Appendix D, we add more insights on the performance of the $T_{bond}$ test statistic
as observed in Table 1.
@@
An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power. This can be observed when comparing the power values of the small-world network to the stochastic block model. This can be attributed to a reduction of sample space of imputed values of the test statistic. Hence, a higher degree of clustering increases the cluster sizes, restricting the direct treatment assignment. For instance, an observed graph in the case of the small-world network had 16 clusters, compared to 11 clusters in the case of the stochastic block model.
@@
We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al. (2018) and Puelz et al. (2022). As discussed in Section 3, we encounter the degenerate null distribution and do not present the corresponding results. This is more prominent in the method proposed by Athey et al. (2018), compared with Puelz et al. (2022). Since Puelz et al. (2022) considers the experiment’s design when selecting focal units, we see more robustness of the biclique test (Puelz et al. (2022)) where degenerate null distributions are reduced in frequency. Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations. We also see that the $T_{bond}$ test statistic performs slightly better than $T_{quant}$ test statistic in some cluster-randomized setups. This can be attributed to $T_{bond}$ being weighted by treated numbers in the graph instead of the corresponding quantile-based weights in $T_{quant}$ and the heterogeneity of the clustered assignment setup. We emphasize that the $T_{bond}$ in conjunction with our methodology continues to be valid regardless of the size of the direct effect. The method proposed by Bond et al. (2012) does not work for settings with non-zero direct effects (Refer Athey et al. (2018) for an elaborate discussion).
@@ 6 Weather insurance adoption in rural China
We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al. (2015)). The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption. The experiment involved intensive training sessions to promote crop insurance products to farmers against extreme weather. A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges. These are important questions to be answered in this setting as they are consequential to government policy action. A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community. An important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
We discuss the experimental design setup by the researchers. Researchers hypothesize that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community. To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance. To capture information dissemination across the social network, these sessions were conducted in two rounds with a gap of 3 days. The idea behind this setup is to allow the farmers some time to discuss the insurance product among themselves. The interference effect is identified by examining whether households in round 2 with more friends who received intensive sessions in round 1 displayed an enhanced insurance take-up rate. Second-round participants were further randomized into three groups. After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not. This additional information randomization of the second group was done to derive the primary information diffusion mechanism. For illustration purposes, we do not analyze this part of the experiment.
@@
7 Conclusion
We develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup. Our proposed method assumes natural distributional conditions in the network-generating process to build a quasi-randomization test. While promising, several open problems remain. Incorporating covariates or nodal attribute information to characterize the random graph null model better may help enhance the power of the test and is an interesting direction for future work. We also assume fully and correctly specified information about the given network. Analyzing the random graph null models under partial information and network misspecification is another exciting avenue for future work. Finally, our focus in this paper has been on randomized experiments exclusively. Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and essential extension is to study random graph null testing procedures in observational studies.
Acknowledgments. We thank Avi Feller, Fabrizia Mealli, Panos Toulis, ACIC 2024 conference participants, and the Tom Ten Have Award committee for their encouragement and feedback. We also thank Sai Sriramya Gorripati, Parshuram Hotkar, Sumit Kunnumkal, and ISB seminar participants for their enthusiasm and comments. We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access.

### v3

@@ 6.2 Test for exponential random graph model
Many sufficient statistics used in real-life social network modeling follow this property. As discussed, some graph statistics include the number of triangles, degree sequence, and number of stars of size two, and they are permutation invariant. We model the data-generating process as an exponential random graph with a permutation invariance property and present a permutation testing procedure.
@@
    $G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\;|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}],G\equiv_{d}G_{obs}\}.$
@@
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a cluster randomized experiment. We assume that $G_{obs}\sim P_{G}$ where $G$ is a permutation-invariant exponential random graph model. Let $G_{X_{i}}=deg(i)$ where $G_{X_{i}}$ denote the graph covariate(s) of unit $i$. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 3 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
@@ 6.3 Test statistic
It should be noted that the above test statistics can only be defined if the sub-populations under consideration are non-empty. For example, if no control unit exists with all neighbors in control or the number of such units is few, we deal with either a definition or a power issue. Thus, we propose a generalization of the above-stated test statistic for robustness. We look at the difference in average treatment effect among control units with a greater than 0.75 quantile of the proportion of treated units and control units with a less than 0.25 quantile of the proportion of treated units. Let $T_{quant_{c}}(G|Z_{obs})=$
$\hskip-42.67912pt\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}$ (7)
@@
We remark that one can potentially improve the test statistic with the knowledge of appropriate covariates or nodal attributes. This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)). One can also build test statistics based on models for interference structures. For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993). We also consider more test statistics, given that the choice is not unique, and different test statistics would differ in their power. We take the test statistic proposed in Bond et al. (2012), which takes the difference between the average over all the edges where the neighbor is a treated or control unit.
@@ 7.1 Data generating mechanism
We now present the data-generating process considered in the simulations. Our potential outcome function is linearly separable into direct treatment effect and spillover effect. The spillover mechanism is a function of the unit’s neighbors’ direct treatment assignment. The potential outcome function also has an additive heterogeneous effect coming from a network-based property, free of a direct treatment assignment vector. We propose the mechanism to be contingent upon the proportion of treated neighbors.
$Y_{i}(z)=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\beta_{\text{deg}}\cdot\frac{\sum_{j=1}^{N}G_{ij}}{\text{max}_{i}{\sum_{j=1}^{N}G_{ij}}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (8)
@@
We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated. Without loss of generality, we take the mean of the normal distribution to be zero, but it can be arbitrarily chosen. We first present the setup for a completely randomized experiment. The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and assigning them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). We state that $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have $599$ total units and $300$ treated units in the model in one case. We study the setup for two generating processes: small-world networks and the stochastic block model. Both models exhibit a high degree of clustering, characterized by the clustering coefficient. The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible. Consider a regular graph where all the nodes have the same cardinality of connections. Let us denote the number of connections a unit has in a regular graph by $K$. Such graphs have a high clustering coefficient, but the distance between two nodes in such graphs is also high. The average distance between two nodes in real-life networks is low (Amaral et al. (2000)). To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$. This parametrization characterizes a small-world network. The higher the rewiring probability, the lower the average path length. Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model. Consider a graph of size $N$ with a partition of nodes. The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$. These parts represent clusters in the graph, and we can model a random graph using them by defining the probability of an edge between nodes within a part and between two parts. Hence, we obtain a probability (symmetric) matrix of size $K\times K$. Note that because we also want a high degree of community structure in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements. The probability matrix is also called the preference matrix in the literature. We also have a coefficient of network dependence in the potential outcome model, denoted by $\beta_{\text{deg}}$. We take $\beta_{\text{deg}}=0$ in Table 2 for a comparative study with the state-of-the-art literature (Athey et al. (2018)). We now define the setup for the simulation study:
@@
    Network: We first present results taking network distribution as a small-world network with the following parameters: K (initial connection of all units) = 10 and $p_{rw}$ (probability of rewiring) = 0.1. We then consider a stochastic block model with block sizes $\{50,100,40,110,299\}$. The preference matrix, which specifies the probability of an edge between two groups, is given below:
@@
The null distribution was generated by sampling 1000 graphs with the same (labelled) degree sequence as the observed graph, and degree-preserving isomorphic to the observed graph in a second option.
Table 1: Rejection rates for the null hypothesis of no spillover effects for a completely randomized experiment
$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$ $T_{elc}$
Small world network 0 0 0.045 0.049 0.051 0.045 0.050
4 0 0.054 0.046 0.046 0.050 0.046
0 0.4 0.187 0.148 0.193 0.163 0.112111We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
4 0.4 0.074 0.151 0.071 0.156 0.060222We obtain the same rejection rate in our replication as reported in Athey et al. (2015).
Stochastic block model 0 0 0.045 0.062 0.047 0.046 0.054
4 0 0.050 0.055 0.053 0.056 0.052
0 0.4 0.191 0.217 0.200 0.206 0.065
4 0.4 0.064 0.220 0.055 0.214 0.048
The p-value is calculated as the sample proportion of times the imputed test statistic for the sampled graph realizations is more than the observed test statistic. We reject the null hypothesis if the p-value obtained is less than the chosen significance level of 0.05. We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null. We present this result, namely rejection rate, in the Table 1. For robustness, we perform the simulations for two network-generating processes. We obtain the rejection rates for the same setup for the procedure proposed by Athey et al. (2018). We implement random focal vertex selection and show our results for the Edge-Level-Contrast test statistic (Athey et al. (2018)). We label this test as ‘Conditional Focal Test’.
We then perform the simulations for a cluster-randomized experimental setup. We present the results in Table 2. The total number of units in the simulation is the same as before, that is, 599. We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Supplementary material A). We take epsilon to be equal to 3. This implies all the clusters contain units less than three hops away from each other, whereas the between-cluster distance is at least three hops for any two units. We take the probability of treating a cluster to be 0.5. We implemented a Bernoulli trial at the cluster level. Under Procedure 3, the null distribution is approximated by taking 1000 samples of graphs isomorphic to the original graph, such that the corresponding subgraphs on treated units and controlled units remain isomorphic and the labelled degree sequence remains preserved. This is obtained by considering permutations where treated units can be mapped only to treated units (thus, control units can only be mapped to control units), and mapping only occurs amongst two units with the same degree. We replicated this setup 4000 times and calculated the proportion of times the proposed test rejected the null. We label Procedure 3 as ‘Block Isomorphism Permutation Test’ in Table 2. This was performed for the two network-generating processes as before.
@@
Small world network 0 0 0.053 0.046
4 0 0.051 0.054
0 0.4 0.420 0.360
4 0.4 0.435 0.360
Stochastic block model 0 0 0.046 0.050
4 0 0.046 0.052
0 0.4 0.197 0.206
4 0.4 0.211 0.214
For robustness, we also perform simulations for the setup where there is a dependence of the potential outcome function on the graph, devoid of direct treatment assignment. For this, we set $\beta_{\text{deg}}=0.4$, installing a positive correlation with the degree of the unit and its corresponding potential outcome. We present the results in Table 3. The simulations are replicated for both the completely randomized experiment and the cluster randomized experiment. The procedures deployed are the degree-preserving isomorphism permutation test and the degree-preserving block isomorphism permutation test. The rest of the setup remains the same as before.
Table 3: Rejection rates for the null hypothesis of no spillover effects for the outcome model including graph covariates. Here, $\beta_{deg}=0.4$ and the Procedures deployed are degree-preserving isomorphism/block isomorphism permutation test for a completely/cluster randomized experiment.
Network Distribution Direct Effect Spillover Effect Completely Randomized Experiment Cluster Randomized Experiment
$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$
Small world network 0 0 0.052 0.045 0.048 0.046
4 0 0.050 0.056 0.054 0.053
0 0.4 0.184 0.162 0.421 0.370
4 0.4 0.065 0.168 0.415 0.330
Stochastic block model 0 0 0.050 0.048 0.050 0.048
4 0 0.051 0.050 0.046 0.043
0 0.4 0.197 0.216 0.210 0.205
4 0.4 0.071 0.208 0.186 0.202
@@ 7.2 Comparison with related work
We see that our proposed methods have type 1 error controlled appropriately. This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1, 2, and 3. When $\tau_{spill}$ equals 0.4, our method performs with competitive power compared to existing methods in the literature (Athey et al. (2018)). We highlight the $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method significantly improves. The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al. (2015)). We see in Table 1 that our proposed method boosts power when the clustering of the network-generating process increases, as with the stochastic block model. This contrasts with the method proposed by Athey et al. (2018), where the method does not retain power. In Supplementary material D, we add more insights on the performance of the $T_{bond}$ test statistic as observed in Table 1.
@@
An interesting result is in the case of cluster-randomized experiments, where a higher degree of clustering shows a drop in power. This can be observed when comparing the power values of the small-world network to the stochastic block model. This may be attributed to a reduction of sample space of imputed values of the test statistic. Hence, a higher degree of clustering increases the cluster sizes, restricting the direct treatment assignment. For instance, an observed graph in the case of the small-world network had 16 clusters, compared to 11 clusters in the case of the stochastic block model.
@@
We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al. (2018) and Puelz et al. (2022). As discussed in Section 3, we encounter the degenerate null distribution and do not present the corresponding results. This is more prominent in the method proposed by Athey et al. (2018), compared with Puelz et al. (2022). Since Puelz et al. (2022) considers the experiment’s design when selecting focal units, we see more robustness of the biclique test (Puelz et al. (2022)) where degenerate null distributions are reduced in frequency. Overall, our method is more robust to degeneracy and performs well in terms of power in our simulations. We also see that the $T_{bond}$ test statistic performs slightly better than the $T_{quant}$ test statistic in some cluster-randomized setups. This may be attributed to $T_{bond}$ being weighted by treated numbers in the graph instead of the corresponding quantile-based weights in $T_{quant}$ and the heterogeneity of the clustered assignment setup. We emphasize that $T_{bond}$ in conjunction with our methodology continues to be valid regardless of the size of the direct effect. The method proposed by Bond et al. (2012) does not work for settings with non-zero direct effects (Refer Athey et al. (2018) for an elaborate discussion). Our method also retains the same power with graphical confoundedness of the potential outcome function, as shown in Table 3 where $\beta_{\text{deg}}$ = 0.4.
@@ 8 Weather insurance adoption in rural China
We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al. (2015)). The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption. The experiment involved intensive training sessions to promote crop insurance products to farmers against extreme weather. A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges. These are important questions to be answered in this setting, as they are consequential to government policy action. A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community. An important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
We discuss the experimental design setup by the researchers. Researchers hypothesize that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community. To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance. To capture information dissemination across the social network, these sessions were conducted in two rounds with a gap of 3 days. The idea behind this setup is to allow the farmers some time to discuss the insurance product among themselves. The interference effect is identified by examining whether households in round 2 with more friends who received intensive sessions in round 1 displayed an enhanced insurance take-up rate. Second-round participants were further randomized into three groups. After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not. The randomization of the second group was done to derive the primary information diffusion mechanism. For illustration purposes, we do not analyze this part of the experiment.
@@
9 Discussion
We develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup. Our proposed method assumes natural distributional conditions in the network-generating process to build a quasi-randomization test. While promising, several open problems remain. Incorporating covariates or nodal attribute information to characterize the random graph null model better may help enhance the power of the test and is an interesting direction for future work. The possibilities of incorporating advanced graph theoretic sampling techniques, and their effect on power, are another interesting line of pursuit. We also assume fully and correctly specified information about the given network. Analyzing the random graph null models under partial information and network misspecification is another exciting avenue for future work. Finally, our focus in this paper has been on randomized experiments exclusively. Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and essential extension is to study random graph null testing procedures in observational studies.
Acknowledgement
We thank Avi Feller, Fabrizia Mealli, Panos Toulis, ACIC 2024 conference participants, and the Tom Ten Have Award committee for their encouragement and feedback. We also thank Sai Sriramya Gorripati, Parshuram Hotkar, Sumit Kunnumkal, and ISB seminar participants for their enthusiasm and comments. We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access. The IMS 2024 ICSDS Student Travel Award partially supported Supriya Tiwari’s research. Pallavi Basu’s research is partially supported by the SERB MATRICS award MTR/2022/000073.
Supplementary material
The Supplementary Material contains proofs, additional simulation results, and algorithmic details.